2x+y=7
help me step by step on this problem please

Answer by sajankumar(7)

solution
2x+y=7
to get the the number for y make x=0
2(0)+y=7
0+y=7
y=7
to get the the number for x make y=0
2x+0=7
2x=7
2x/2=7/2 divide both sides with 2
x=7/2